# Seeing Clearly, Answering Incorrectly: A Multimodal Robustness Benchmark for Evaluating MLLMs on Leading Questions

### v1

@@ B.3 What constitutes an effective pair of positive and negative samples for training?
Appendix D Licensing and Hosting
Author Statement.
We bear all responsibilities for the licensing, distribution, and maintenance of our dataset.
License.
MMR benchmark and MMR-data are under CC BY 4.0 license.
Hosting.
MMR benchmark and MMR-data can be viewed and downloaded on GitHub at https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark or on Huggingface at https://huggingface.co/datasets/BAAI/Multimodal-Robustness-Benchmark. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in JSON file format.
Metadata.
Appendix E Datasheet and Maintenance Plan
E.1 Motivation
For what purpose was the dataset created?
Answer: The MMR benchmark is designed to evaluate MLLMs’ comprehension of visual content and robustness against misleading questions, ensuring models truly leverage multimodal inputs rather than relying solely on textual reasoning. The MMR-data is generated to enhance MLLMs’ understanding capability and robustness by providing a training set with paired positive and negative visual question-answer samples.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
Answer: The MMR benchmark and MMR-data are created by Yexin Liu (BAAI, HKUST(GZ)), Zhengyang Liang (BAAI, BUPT), Yueze Wang (BAAI), Muyang He (PKU), Jian Li (Youtu Lab, Tencent), Bo Zhao (BAAI)
Who funded the creation of the dataset?
Answer: 1) Beijing Academy of Artificial Intelligence. 2) NSFC-62306046.
E.2 Composition
What do the instances that comprise the dataset represent? (e.g., documents, photos, people, countries)
Answer: Each instance in the MMR benchmark represents an image, a question, 4 options, and a standard answer. Each instance in the MMR-data includes an image, a question, and an answer. Images are stored in JPG or PNG file format, while the questions, options, and standard answers are all stored in JSON format files.
How many instances are there in total (of each type, if appropriate)?
Answer: For the MMR benchmark, we collect 600 instances, consisting of 300 positive and 300 negative questions. The specific distribution of question types is detailed in Fig. 4 in the appendix. For MMR-data, we generate 112k samples (56k positive and 56k negative) using GPT-4o, with an average of 6.82 rounds of conversation per sample.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
Answer: All instances in the MMR benchmark and MMR-data were newly annotated by us or generated by GPT-4o.
Is there a label or target associated with each instance?
Answer: Yes, each instance provides the correct answer with an explanation.
Is any information missing from individual instances?
Answer: All instances are complete.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
Answer: For the MMR benchmark, after determining the question type, we then look for diverse instances. Each image is independent and not explicitly connected to others and two questions (positive and leading questions) are annotated with wwith each image. For MMR-data, some instances may have the same image but different questions and answers. Each instance clearly indicates the corresponding image, and each image has a unique identifier.
Are there recommended data splits (e.g., training, development/validation, testing)?
Answer: Yes, the MMR benchmark is specifically designed for evaluation, while the MMR-data is generated for training.
Are there any errors, sources of noise, or redundancies in the dataset?
Answer: No.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
Answer: The annotations of the MMR benchmark are manually annotated by us, whereas the annotations of MMR-data are generated by GPT-4o. All metadata will be publicly accessible in the dataset repository. The images are collected from websites or open-source datasets. We have taken measures such as changing the resolution, modifying the aspect ratio, changing the format, and editing to minimize the impact on the original work rights. However, it is still possible that the copyright holder may request the deletion of certain data. If this happens, we will edit the content without affecting the questions and answers. If retaining images is also not allowed, we will still keep the annotation data and provide metadata (including URL) for the corresponding images.
Does the dataset contain data that might be considered confidential?
Answer: No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
Answer: No.
E.3 Collection Process
The data collection process for the MMR benchmark is detailed in Appendix A. 2 of the main paper. When labeling the MMR benchmark, those involved in the labeling search for suitable images based on pre-defined question types, type definitions, questions, and answer requirements, and then label them. The data collection process for MMR-data is outlined in Sec. 4 of the main paper.
E.4 Uses
Has the dataset been used for any tasks already?
Answer: Yes, the MMR benchmark has been used to evaluate the understanding capabilities and robustness to leading questions in 18 different MLLMs. The MMR-data has been used to train three different MLLMs, validating the effectiveness of the dataset.
What (other) tasks could the dataset be used for?
Answer: The MMR benchmark is primarily used to comprehensively evaluate the understanding capabilities and robustness to leading questions in MLLMs. Given its inclusion of multiple different tasks, it can also be used to individually assess the ability of MLLMs on specific tasks, such as image-grounded reasoning.
Is there a repository that links to any or all papers or systems that use the dataset?
Answer: No.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
Answer: The annotations of the MMR benchmark are manually annotated by us, whereas the annotations of MMR-data are generated by GPT-4o. All metadata will be publicly accessible in the dataset repository. The images are collected from websites or open-source datasets. We have taken measures such as changing the resolution, modifying the aspect ratio, changing the format, and editing to minimize the impact on the original work rights. However, it is still possible that the copyright holder may request the deletion of certain data. If this happens, we will edit the content without affecting the questions and answers. If retaining images is also not allowed, we will still keep the annotation data and provide metadata (including URL) for the corresponding images.
Are there tasks for which the dataset should not be used?
Answer: The MMR benchmark and MMR-data cannot be used to evaluate the image or video generation capabilities of the MLLM.
E.5 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Answer: Yes. The MMR benchmark and MMR-data are publicly available on the Internet.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
Answer: The MMR benchmark and MMR-data are available on GitHub https://github.com/BAAI-DCAI/Multimodal-Robustness-Benchmark or on Huggingface at https://huggingface.co/datasets/BAAI/Multimodal-Robustness-Benchmark.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
Answer: CC BY 4.0.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
Answer: No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
Answer: No.
E.6 Maintenance
Who will be supporting/hosting/maintaining the dataset?
Answer: The authors will be supporting, hosting, and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Answer: Please contact one of the authors (yliu292@connect.hkust-gz.edu.cn; 1ce@bupt.edu.cn; bozhaonanjing@gmail.com)
Is there an erratum?
Answer: No. We will make announcements if there are any.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
Answer: People may appear in the reference images. People may contact us to exclude specific data instances if they appear in the reference images.
Will older versions of the dataset continue to be supported/hosted/maintained?
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
Answer: If others wish to add data, they can apply to do so provided the data is compliant and reasonable. However, making other modifications based on our dataset is currently not allowed.

### v2

@@ Appendix E Examples of Paired Positive and Negative Samples Generated
Appendix F The Details of “Instruction prompt”.
STEP-1 (DO NOT OUTPUT): Image Information Extraction 1. Text/Numbers: Identify clear text or numbers in the image and their positions. 2. Objects: Identify objects, their properties (name/category/color/texture/shape/pose), position, and quantity. 3. People: Identify people, and their attributes (name, position, gender, facial expression), and distinguish between real people and portraits in photos or posters. 4. Object Relationships: Understand relationships between objects (relative positions or hierarchy). 5. People-Object Relationships: Understand relationships between people and objects (e.g., someone riding a bike). 6. Activities: If there are multiple people, identify their activities. 7. Events: Understand events in the image and their cultural, emotional, or knowledge-related context. STEP-2 (DO NOT OUTPUT): Carefully review the image and the question to ensure an accurate understanding of each question. Example Responses: Question: Is there a red bicycle in the image?
Answer: Yes, there is a red bicycle in the image. It is clearly shown in the center.
Question: Is the car in the image green?
Answer: No, the car in the image is blue. The color is distinct.
STEP-3: Answer with the option’s letter from the given choices directly.
Figure 17: Common Prompt for information extraction.
Appendix G Details about Content Guided Refinement Strategy
G.1 Prompt for extraction visual information
The prompt is shown in Fig. 18.
Extract information from the image and structure the output as follows: 1. Text/Numbers: Identify any visible text or numbers, including their content, font, and precise positions in the image. 2. Objects: Identify each object, detailing its name/category, color, texture, shape, pose, relative location within the image, and quantity. 3. People: Identify all visible individuals, with attributes such as position, gender, facial expression, and distinguish between real individuals and representations (e.g., in portraits, posters). 4. Relationships & Interactions: Describe spatial or contextual relationships between objects, and interactions or spatial relationships between people and objects. 5. Activities: Identify and describe any observed activities or actions involving people or objects. 6. Events: Identify any significant events or scenes occurring in the image.
Figure 18: Prompt for information extraction.
G.2 Examples of Content Guided Refinement Strategy Results
The example is shown in Fig. 19.
Refer to caption Example ID: attribute/existence/COCO_val2014_000000170365
Image: COCO_val2014_000000170365.jpg
Question: What is the object under a plate of yellow plants in the picture?
Options:
A. Towel
B. Shower enclosure
C. Sink
D. Closestool
Ground Truth: D
Question Type: POS
Info:
1. Text/Numbers: There is no visible text or numbers in the image. 2. Objects: (a) Sink: A white porcelain sink with a silver faucet. (b) Toilet: A white porcelain toilet. (c) Bathtub: A white porcelain bathtub with a shower door. (d) Mirror: A framed mirror on the wall above the sink. (e) Potted Plant: A yellow plant in a pot on the counter. (f) Towels: White towels hanging on the towel rack. (g) Toilet Paper: A roll of toilet paper on the back of the toilet. (h) Tissue Box: A box of tissues on the counter. (i) Trash Can: A white trash can with a lid. (j) Toilet Brush: A white toilet brush in a holder. 3. Relationships & Interactions: Various spatial and functional relationships between objects, as described above. 4. Activities: None observed. 5. Events: The image shows a typical bathroom setup.
Figure 19: Examples of Content Guided Refinement Strategy Results.
Appendix H Details about Visual Attention Refinement Strategy
H.1 Detailed Description
Our analysis indicates that the probability of the model generating correct answers decreases when the attention values between the question token and the visual token are relatively low. To address this, we propose leveraging the attention mechanism between the question and visual tokens to optimize the visual prompts, as shown in the algorithm  1. Specifically, we extract the attention parameters from the final transformer layer of the MLLM, while other proposed steps remain consistent with those outlined in the main text. Using these parameters, an attention map is generated, where regions with higher values indicate areas of focus for the model. To guide the model toward regions relevant to the question, we invert the attention map, apply regularization and filtering, and blend it with the original image. The blending process employs coefficients of 0.85 for the original image and 0.15 for the attention map.
Algorithm 1 Optimizing Visual Prompts Using Attention Mechanisms
0:   MLLM with pre-trained parameters, input question $Q$, input image $I$, blending coefficients $\alpha=0.85$, $\beta=0.15$
0:   Optimized visual prompt $I^{\prime}$ for inference
1:   Extract attention parameters $A$ from the final transformer layer of the MLLM.
2:   Compute the attention map $M$ using $A$ between the question token and the visual tokens.
3:   Identify high-value regions in $M$ where the model focuses during inference.
4:   Invert the attention map:
5:   Compute the inverted attention map $M_{\text{inv}}=1-M$.
6:   Regularize and filter:
7:   Apply normalization and filtering to $M_{\text{inv}}$ to obtain $M_{\text{filtered}}$.
8:   Blend attention map with the original image:
9:   Generate the optimized image $I^{\prime}$ as:
$I^{\prime}=\alpha\cdot I+\beta\cdot M_{\text{filtered}}$
10:   return  Optimized image $I^{\prime}$ for inference.
H.2 Examples of visual attention refinement strategy
Refer to captionRefer to caption POS Example ID: attribute/existence/21
Image: 21.jpg
Question: How many puppies are there in the picture?
Options:
A. There is only one puppy in the picture and the other one is a cat.
B. There are two puppies in the picture.
C. There are no puppies in the picture, just two kittens.
D. There is no puppy in the picture, just two stuffed animals.
Ground Truth: A
Question Type: POS
Question: Does the picture show a dog and its pups?
Options:
A. The pup lay on top of the dog.
B. The picture shows a dog and its two puppies.
C. There are no puppies in the picture, just two kittens.
D. There is a dog in the picture and the other smaller animal is a cat.
Ground Truth: D
Question Type: NEG
Info:
1. Text/Numbers: There is no visible text or numbers in the image. 2. Objects: (a) Dog: White and brown, smooth fur, oval body, sitting on a blanket. (b) Cat: White and gray, soft fur, pointed ears, sitting on a blanket. (c) Blanket: White, soft, fluffy, laying on a wooden surface. (d) Wooden Surface: White, rough and uneven, under the blanket. 3. People: No visible people. 4. Identification: No real individuals, likely a portrait or poster. 5. Activities: No observed actions. 6. Events: No significant events.
Figure 20: Examples of visual attention refinement strategy.
The example is shown in Fig. 20.
Appendix I Broader Impact and Limitation
Broader impact.
Limitation.